A savings account at Xyz Bank offers 4% interest, compounded annually. How much money must I invest today in order for the account to grow to 5000$ in ten years? Answer to the nearest whole dollar.
Answer by jorel555(1290)

Since it compounds annually, your annual yield is 100% + 4%, or 1.04 times your initial investment. If you want to know how much that is in ten years, then:
I * (1.04)^10= the value of your account after ten years. So
I*1.04^10=5000
I=5000/1.04^10=5000/1.48=3378.38 original investment needed to have 5000 at the end of ten years!!!!!!!